11. What is the molarity of a solution containing 15.4 mg of NaCl in 0.5 ml of dH,O?
Answer #1

Molarity is the number of moles of solute per litre of solution.

No. Of moles of NaCl = weight of NaCl/ molecular weight of NaCl

Molecular weight of NaCl = 58.4g

Weight of NaCl = 15.4 mg = 0.0154 g

No.of moles of NaCl = 0.0154/58.4 = 0.0002635

Molarity = 0.0002635/ 0.0005 = 0.527 M